Could I get some help on these two please?
"Factor the quadratic expression completely, and find the roots of the expression"
a. 135x^2 - 222X + 91
b. 6x^2 - 42x + 72
Thank you very much!

Answer by checkley71(8403)

135X^2-222X+91=0
USING THE QUADRATIC EQUATION X=(-B+-SQRT[B^2-4AC])/2A WE GET
X=(222+-SQRT[222^222-4*135*91])/2A
X=(222+-SQRT[49,284-49,140])/270
X=(222+-SQRT[144])/270
X=(222+-12)/270
X=210/270
X=.777777
X=234/270
X=.866666
6X^2-42X+72=0
X^2-7X+12=0
(X-4)(X-3)=0
X-4=0
X=4
X-3=0
X=3